OpenStudy (tennis5518):

the distance between spice island and kay island is 3.4 in on the map if the scale on the map is 1 in=25 miles what is the actual distance between the two island

OpenStudy (tennis5518):

Mathlete @freckles

Mehek (mehek14):

make proportions \(\dfrac{1}{25}=\dfrac{3.4}{x}\)

OpenStudy (tennis5518):

oh ok

Mehek (mehek14):

cross multiply 25 * 3.4 = 1 * x

OpenStudy (tennis5518):

85

OpenStudy (tennis5518):

@mehek14

OpenStudy (skullpatrol):

Don't forget the units :-)

OpenStudy (tennis5518):

oh yea

Mehek (mehek14):

so that's the distance

OpenStudy (tennis5518):

85miles

OpenStudy (tennis5518):

Thxs both!! :)

OpenStudy (skullpatrol):

Thanks for asking :)

Mehek (mehek14):

yw :)